OpenStudy (anonymous):

Given the function g(x) = 4x - 5, compare and contrast g(2) and g(-4). Choose the statement that is true concerning these two values.

hero (hero):

g(x) = 4x - 5 g(2) = 4(2) - 5 g(2) = 8 - 5 g(2) = 3 g(x) = 4x - 5 g(-4) = 4(-4) - 5 g(-4) = -16 - 5 g(-4) = -21 I would choose the statement that is true concerning these two values, but you haven't posted it yet.